### v2

# Aesthetics is Cheap, Show me the Text: An Empirical Evaluation of State-of-the-Art Generative Models for OCR

Abstract
Text image is a unique and crucial information medium that integrates visual aesthetics and linguistic semantics in modern e-society. Due to their subtlety and complexity, the generation of text images represents a challenging and evolving frontier in the image generation field. The recent surge of specialized image generators (e.g., Flux-series) and unified generative models (e.g., GPT-4o), which demonstrate exceptional fidelity, raises a natural question: can they master the intricacies of text image generation and editing? Motivated by this, we assess current state-of-the-art generative models’ capabilities in terms of text image generation and editing. We incorporate various typical optical character recognition (OCR) tasks into our evaluation and broaden the concept of text-based generation tasks into OCR generative tasks. We select 33 representative tasks and categorize them into five categories: document, handwritten text, scene text, artistic text, and complex & layout-rich text. For comprehensive evaluation, we examine six models across both closed-source and open-source domains, using tailored, high-quality image inputs and prompts. Through this evaluation, we draw crucial observations and identify the weaknesses of current generative models for OCR tasks. We argue that photorealistic text image generation and editing should be internalized as foundational skills into general-domain generative models, rather than being delegated to specialized solutions, and we hope this empirical analysis can provide valuable insights for the community to achieve this goal. This evaluation is online and will be continuously updated at our GitHub repository.
[…]
1 Introduction
Generating images with machines represents humanity’s ambitious pursuit to translate the visual world into algorithms, teaching computers the fundamental human skill of creation. Since the 1960s, researchers have started to use computers for creating artworks like films and still ASCII art [1, 2, 3, 4]. Later, this field evolved from early human-programmed, procedural rules [5, 6] to modern, data-driven artificial intelligence (AI) approaches, exemplified by the early Variational AutoEncoders (VAE) [7, 8] and Generative Adversarial Networks (GAN) [9, 10]. Beyond image-to-image translation or conditional image generation [11], researchers sought to control the generation process more naturally and flexibly. This fueled the emergence of text-to-image (T2I) generation [12], allowing users, even those non-technical, to generate visual content with natural language. Building upon the proliferation of T2I generation [13, 14, 15], the community has now broaden the application realm into instruction-guided image generation and editing, driven by the confluence of large-scale text-image datasets, advances of language models [16, 17], and innovative architectures like diffusion [18, 19], flow matching [20], and autoregressive generation [21, 22].
Mainstream state-of-the-art (SOTA) generation approaches fall into two types: specialized image generation models and native unified generative models. The first type is primarily built based on diffusion or flow matching framework, such as DALL·E [23, 24, 25], Imagen [26], Stable Diffsuion [19, 27, 28], Flux [29, 30], and OminGen [31, 32]. They focus merely on generating or editing images based on instructions, using a relatively small text encoder [16] for textual prompt encoding. In contrast, the second type unifies image understanding and generation within one Transformer model, termed unified generative models. They are typically built based on multimodal large language models (MLLMs) with the autoregressive objective, which can be further divided into three meta-categories. (1) Models w/o diffusion modeling. This line of models use VQGAN [33] or VQVAE [34] to tokenize images into discrete tokens as input to the language models [35, 36] for multimodal autoregressive learning, and detokenizes the output features back into images [37, 38, 39, 40, 41, 42, 43, 44]. (2) Diffusion-embedded models. In contrast, this type of works embrace diffusion modeling, integrating diffusion’s loss and the iterative denosing process into the Transformer for training and inference, respectively [45, 46, 47, 48, 49, 50]. They can switch their role as language models for autoregressive text generation or a diffusion model for parallel image synthesis. (3) Models with extra diffusion head. Compared to the second meta-category, this fashion of approach retains only autoregressive next-token prediction loss for training, while setting up an additional, lightweight diffusion head [22] alongside the MLLM, which is conditioned on the MLLM’s hidden states for image generation [51, 52, 53, 54, 55]. Recently, the release of Gemini-2.0-Flash [56] and GPT-4o [57] marks as a milestone of unified native image generation, exhibiting commercial-level image generation and editing capabilities with non-deteriorated understanding skills.
Figure 1: Task categorization and subordination. We classify 33 OCR generative tasks into five primary categories based on text characteristics: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text. Each primary category encompasses multiple sub-tasks.
Text images, such as documents, scene texts, and handwriting, represent special and critical visual information carriers for human communication. Unlike natural objects or paintings, text images contain structured linguistic information that requires both visual perception and language understanding for proper interpretation. This dual requirement makes the generation and manipulation of text images a particularly challenging frontier in the generative AI community [58, 59]. Despite significant advances in large-scale T2I synthesis, generating high-quality text images remains problematic due to fundamental challenges, such as ensuring precise character formation and legibility across scales, maintaining semantic consistency between text and visual content, and handling diverse layouts and orientations. General-domain generators excel at natural image generation but usually falter in text image generation and editing, and very few of them [28, 57, 60] have formally acknowledged or addressed this critical deficiency. We argue that photorealistic text generation is an important and indispensable ability for achieving artificial general intelligence (AGI). Although numerous models tailored for text image synthesis have been proposed [61, 62, 63, 64, 65, 66], text generation should be internalized as a foundational skill, rather than delegated to specialized solutions. In addition, a comprehensive assessment dedicated to text image generation/editing is lacking. Such assessment is essential for systematically evaluating the strengths and weaknesses of existing models, illuminating clear pathways for future research.
To this end, we present a comprehensive evaluation of text image generation and editing capabilities in SOTA generative models. In literature, the automatic interpretation of text images is known as optical character recognition (OCR) [67], among which many tasks are inherently generative tasks, such as document deshadowing [68], scene text removal [69], and font generation [65]. Hence, we broaden the definition of text generation tasks and reframe them as OCR generative tasks. Our evaluation operates in three stages. First, we curate 33 OCR tasks and classify them into five main categories based on text properties: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text. Second, we selected a representative suite of models for benchmarking, including both specialized generators and unified models from closed-source and open-source domains: GPT-4o [57], Qwen-VLo-Preview [60], Flux.1-Kontext-dev [30], OmniGen2 [31], BAGEL [50], and Janus-4o [44]. Third, we tailor expressive prompts and select high-quality input images from public datasets [70, 71] or web sources for each task, encompassing bilingual evaluation in both English and Chinese. From the evaluation, we discover that text image generation and editing remain substantial challenges for existing generative models, and draw several crucial observations as follows.
  1. 1.
    Awesome creativity. Existing models excel at generating creative and design-oriented images with text, especially posters, slides, and street scenes, when given detailed prompts. Regardless of the text quality, the style, color harmony, and overall composition often demonstrate remarkable artistic appeal and sophistication.
  2. 2.
    Insufficient perception and localization for specific text. In text removal and editing tasks, current models struggle to accurately modify designated regions, often resulting in incomplete modifications or unintended alterations in other areas. This prohibits them from performing fine-grained text content manipulation.
  3. 3.
    Poor structural preservation. While some tasks require modifications to specific text regions, current models often fail to preserve other areas that should remain unchanged. For example, in text removal, editing, and layout-aware text generation, models may successfully manipulate the target text but inadvertently alter the style, position, or content of surrounding text and background details. Similarly, document enhancement operations such as dewarping, deshadowing, and deblurring frequently introduce unwanted changes to document layout and content beyond the intended corrections.
  4. 4.
    Unstable instruction-following ability and hallucination. Models occasionally fail to follow user instructions and produce unstable outputs. Flux.1-Kontext-dev frequently generates entirely unrelated images in T2I tasks. Qwen-VLo-Preview (document dewarping), OmniGen2 (document dewarping, document deblurring, and modern/historical document editing), and BAGEL (appearance enhancement) occasionally produce anomalous outputs that deviate from user instructions.
  5. 5.
    Problematic processing of complex content. Existing models typically struggle with complex visual content. In document dewarping, embedded graphics cannot be properly restored. In document deshadowing, tables with hierarchical text and sophisticated structures can be accurately replicated. Also, while they can generate simplified Chinese characters, they cannot handle complex Chinese characters.
  6. 6.
    Suboptimal synthesis of dense, long text images. As the required amount and density of generated text or the text inside the input image increase, the results become more prone to errors. In document-related tasks, the generated text is sometimes blurred due to high density. Paragraph-level handwriting generation showcases more errors compared to line-level results. This phenomenon is further exacerbated given complex fonts, as exemplified in generating long text with artistic font styles.
  7. 7.
    Constrained resolution and aspect ratio control. Outputting correctly edited images while maintaining the original resolution or aspect ratio remains challenging. While GPT-4o excels at image editing (except handwritten images), it outputs dimensions in 512-pixel multiples, inevitably distorting image layouts. Other models like BAGEL and OmniGen2 preserve image aspect ratios but demonstrate weaker editing capabilities.
  8. 8.
    Deficient unified understanding and generation skill. Unified generative models, by virtue of training on visual understanding tasks, are expected to develop better comprehension skills. Yet, except for GPT-4o, the other three models exhibit limited world knowledge to fully and precisely interpret OCR-related instructions, particularly evident in professional tasks like dewarping, deblurring, and historical document restoration. Larger scale of training data for OCR-based analysis and text-centric feature input may alleviate this issue.
  9. 9.
    To-be-improved vertical-domain task performance. For very specialized vertical-domain tasks, such as historical document restoration, historical image style transfer, and modern document tasks, these general-domain models demonstrate inferior proficiency. Incorporating these tasks into model training could improve both general and specialized generation capabilities.
  10. 10.
    Limited multilingual proficiency. While current models are capable of English text generation, some of them, particularly open-source ones, fail to generate Chinese text. The resulting text is often garbled and illegible, with frequent character omissions. Additionally, Chinese text specified in instructions is sometimes incorrectly converted into other languages. Adding multilingual support is essential for universally and linguistically applicable generative models.
  11. 11.
    Large gap between open-source and closed-source models. In terms of text generation and editing, closed-source models, especially GPT-4o, significantly outperform the open-source ones. While this could be limited by the smaller size of open-source models (only 7B), more future efforts should be devoted to improving their text synthesis skills. Scaling laws specifically for text image generation have not been unveiled, representing a promising research direction.
While some prior works [72, 73] have evaluated GPT-4o’s native generation abilities in general object domains, we transition the focus to text images and expand our evaluation to both open-source and closed-source unified generative models. We hope this work provides valuable insights into the strengths and limitations of current generative models, benefiting the design of future models toward better text synthesis abilities.
2 Evaluation Setting
Evaluated tasks. We curate 33 common OCR tasks that can be formulated as generation or editing objectives, including document dewarping, handwritten text removal, scene text editing etc. These tasks are categorized into five main categories: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text, where each main category contains multiple sub-tasks. For each sub-task, we design corresponding English and Chinese prompts as consistent inputs across all tested models. A mindmap introducing detailed task subordination is shown in Fig. 1.
Tested models. Our evaluation considers both closed-source models and open-source models. Closed-source models include GPT-4o [57] and Qwen-VLo-Preview [60]; while open-source models include Flux.1-Kontext-dev [30], OmniGen2 [31], BAGEL [50], and Janus-4o [44]. Except for Flux.1-Kontext-dev and OmniGen2, which are specialized image generation models, all others are unified generative models that are capable of both understanding and generation. In addition, all models support text-to-image (T2I) generation and instruction-based image editing, competent for our evaluation tasks. An overview of these models is summarized in Table 1.
Evaluation. We present qualitative results through visualization for each task. Quantitative metric computation is temporarily not considered. Each visualization presents the used prompts, input images (if any), and output images from all models in a grid format for intuitive comparisons.
Table 1: Overview of tested models. Pub. denotes publication. Und. denotes understanding. Gen. denotes Generation. GPU Memory specifically refers to GPU memory requirements for model inference.
Model Pub. Date Type Availability #Parameters Max Context Length GPU Memory
GPT-4o [57] 2025.3.25 Unified Und. & Gen. Closed-Source - - -
Qwen-VLo-Preview [60] 2025.6.26 Unified Und. & Gen. Closed-Source - - -
Flux.1-Kontext-dev [30] 2025.6.17 Specialized Gen. Open-Source 12B 77 tokens 48GB
OmniGen2 [31] 2025.6.23 Specialized Gen. Open-Source 7B 128,000 tokens 48GB
BAGEL [50] 2025.5.20 Unified Und. & Gen. Open-Source 7B 32,768 tokens 60GB
Janus-4o [44] 2025.6.22 Unified Und. & Gen. Open-Source 7B 4,096 tokens 48GB
3 Result of Document Image
3.1 Modern Document
Figure 2: Results of the document dewarping task. The red glowing indicates the relatively best output corresponding to each input. The overall dewarping results are inferior, where GPT-4o rectifies the image to be flat but loses some embedded graphics and text, while other methods even fail to perform dewarping and lose substantial textual details.
Modern documents plays an essential role in human daily activities for information transmission, consisting of text, graphics, tables, and other informational elements. Therefore, document digitization and processing has become increasingly crucial for efficient information management and automated analysis, with OCR as a core technique [74, 75]. Typically, before performing OCR, the digitized documents should undergo preprocessing to enhance OCR accuracy, such as flattening the scrumpling of the documents, removing shadows, and sharpening the text. These preprocessing technologies constitute different document processing tasks, and most of them can be formulated in an image-to-image translation manner. Given the recent trend of instruction-based image generation and editing, we attempt to investigate whether these models can address important and practical document-related tasks.
Figure 3: Results of the document deshadowing and document deblurring tasks. Flux.1-Kontext-dev exhibits the optimal deshadowing results in both language cases, with most texts preserved and shadow removed. Other models either fail to remove the shadow or mistakenly repeat the textual content. For document deblurring, Flux.1-Kontext-dev showcase nearly perfect result in the English scenario. GPT-4o performs better in the Chinese case with precise text restoration, but fails to restore the document structure.
3.1.1 Document Dewarping
Document dewarping refers to correcting geometric distortions of document images, typically caused by curved, warped, or folded pages when scanned or photographed. Models are required to output a flat document surface with original text preserved, facilitating reading or further OCR processing [76, 77, 78, 79, 80]. Existing dewarping methods compute a displacement field that maps the input, warped image to the output, dewarped image. Yet, from the perspective of image generation, it can be formulated as an image-to-image translation task. Early generative models can not interpret the “dewarping” or “rectification” instructions and handle the drastic and detailed pixel changes to derive an accurate, flat document. Given the advancement of image generators, especially the unified generative models in both understanding and generation, they may now be able to understand the “dewarping” instruction and conduct sophisticated pixel transformation. More broadly, we think this task somehow represents the perception and comprehension abilities of models to the real physics world and therefore conduct a detailed evaluation on this task.
Results of the document dewarping task from different models are presented in Fig. 2. As observed, all the dewarped results are unsatisfactory. GPT-4o is the only model that identifies the instruction of “dewarping” and delivers flattened output documents, showcasing impressive instruction-following capability. While this is a surprising discovery, GPT-4o loses a large portion of original text content and some embedded graphics, also fails to render Chinese content, making the dewarped output infeasible for real-world usage. For other models, they mostly fall short in understanding and following the “dewarping” or “crop” instruction, and generate blurred, chaotic text. Qwen-VLo-Preview even exhibits hallucination without following the input and instructions, as shown in the penultimate row, third column of Fig. 2.
Figure 4: Results of the document appearance enhancement task. Only GPT-4o and Flux.1-Kontext-dev are capable of comprehending the instruction of outputting “PDF-like” documents. However, most of them fall short in preserving document structures, particularly demonstrated by the mistaken repetition of table contents.
Figure 5: Text editing results for modern document. The first three models can follow the instruction (at least partially) while the last three models can not, potentially due to their smaller parameter scale (only 7B). This suggests that document text editing can be addressed by existing generative models when sufficient model scale is available.
3.1.2 Document Deshadowing and Document Deblurring
Document deshadowing refers to removing potential shadow within the document images [81]. Document deblurring refers to removing blur artifacts from document images to restore sharp, readable text and clear visual content [82]. Similar to document dewarping, we require the model to output a deshadowed or a deblurred image according to the input image and instruction. The results are presented in Fig. 3. Flux.1-Kontext-dev demonstrates optimal performance in the deshadowing task, which not only removes the shadow but also preserves the original text content and background color. Other models either fail to remove the shadow (OmniGen2, BAGEL), accidentally repeat text content, or change the document’s background color (GPT-4o, Qwen-VLo-Preview). For deblurring, the English case is easier than the Chinese case. Flux.1-Kontext-dev delivers nearly perfect results with clear text and correct content positions, while GPT-4o seems to automatically complete the whole content and generate a new image, possibly witnessing this document’s content during training. OmniGen2 might misunderstand the instruction as erasing all text, thereby delivering a blank image. In the Chinese case, all models can not deblur the text and keep the original content structure (title, subtitle, main text). Surprisingly, although not a unified understanding and generation model, Flux.1-Kontext-dev demonstrates notable low-level visual document processing abilities, showcasing potential for further development.
3.1.3 Document Appearance Enhancement
Document appearance enhancement, also known as document illumination rectification [83, 84], denotes the technique of mainly correcting uneven lighting conditions and removing other real-world degradations like shadows and bleed-through to improve camera-captured documents’ visual quality. Evaluation results are presented in Fig. 4. In the first row, while Flux.1-Kontext-dev achieves relatively good enhancement and text preservation, it adds unexpected woodgrain edges. BAGEL fails to correct the lighting and outputs red-green-mixed background. Qwen-VLo-Preview and Janus-4o produce chaotic, unreadable text. GPT-4o and OmniGen2 exhibit erroneous text repetition similar to Sec.3.1.2. In the second row, no model can perfectly preserve the table content, and the restoration of Chinese text is unsatisfactory. Furthermore, only GPT-4o and Flux.1-Kontext-dev comprehend the instruction to output “PDF-like” documents, suggesting the need to improve generative models’ understanding of professional document terminologies.
3.1.4 Text Editing
We also evaluate the text editing capabilities of existing models on PDF or real-world documents, with results shown in Fig. 5. We discover that GPT-4o, Qwen-VLo-Preview, and Flux.1-Kontext-dev can follow the instruction to perform target modifications, although the modifications are sometimes incomplete or incorrect. However, the last three models consistently fail to edit documents. This may be attributed to their smaller size, i.e., 7B parameters, whereas the first three models possess larger size (The scales of GPT-4o and Qwen-VLo-Preview are unknown but are likely larger than 7B; Flux.1-Kontext-dev has 12B parameters). This indicates that, for text editing on documents, a larger model size is a fundamental requirement. Still, based on a sufficient model scale, further optimization is needed for multiple editing entries and long-sentence editing.
3.2 Historical Document
Figure 6: T2I generation results for historical documents, which typically feature rich and dense text. GPT-4o fulfills generation requirements in most cases with precise English and Chinese text, despite sometimes the content is incomplete or incorrect. However, other models can at most generate ancient book pages, but fail to generate readable textual content. For complete prompts, please refer to our GitHub repository.
Historical documents represent invaluable repositories of human cultural heritage. In recent years, the digitization and analysis of historical documents [85, 86, 87] have emerged as a critical research area in the community, aiming to aid in the preservation and understanding of these ancient cultural artifacts. These documents contain dense textual content, complex layouts, and antiquated fonts, which present great challenge to image generation models. Therefore, we benchmark existing models on various historical document-related tasks, including general tasks like T2I generation and text editing, as well as vertical-domain tasks like historical document restoration.
Figure 7: Text editing and document restoration results for historical documents. For editing, all models fail to modify specific Chinese text. The relatively best GPT-4o suffers from repeated and wrong editing, while the worst Janus-4o can not even output text but unreadable characters. Yet, some models like GPT-4o and Flux.1-Kontext-dev demonstrate promising edit results on English text editing. For historical document restoration, all models fail to generate new text with the original writing style preserved and keep a consistent background.
3.2.1 T2I Generation
We present the T2I generation result of historical documents in Fig. 6. We design long and informative prompts that describe the text content, font style, writing paper details, and other environmental factors like desks and the oil lamp. We ask the model to generate one-page historical scripts in the first to third rows and a three-page script in the fourth row. As observed, only Flux.1-Kontext-dev successfully generates an authentic ancient book in the first row, whereas all other models produce only script-like images reminiscent of ancient writing. This indicates that Flux.1-Kontext-dev may lack historical document training data. However, while models can somehow generate English text, they consistently fail to render Chinese text. GPT-4o is the only model capable of generating both English and Chinese fonts, and also follows the page requirement in the instructions (one-page or three-page). Other models could fail to follow instructions and produce two pages given the one-page requirement. Still, GPT-4o can produce incomplete or incorrect text sometimes.
Figure 8: Results of the style transfer and super resolution tasks for historical documents. For style transfer, while Qwen-VL-Preview produces the relatively best result by preserving the column structure from image 1 and text sparsity from image 2, the rendered text is blurred and lacks semantics. Other models can not follow the instructions and even produce unreadable results. In the super resolution task, GPT-4o and Qwen-VLo-Preview identify the instruction of improving image resolution, while the others can not understand or misunderstand it. Although the resolution is improved, it comes at the cost of color change or font distortion.
3.2.2 Text Editing and Historical Document Restoration
We then perform text editing on historical documents, with results shown in the top frame of Fig. 7. For Chinese text editing, tested models consistently fail to modify Chinese text images. Although GPT-4o locates and modifies the target text, the edited content is incorrect and nearby text is accidentally altered. For English text, GPT-4o and Qwen-VLo-Preview successfully render the target text “COVERING”, but only GPT-4o places it in the correct place. Other models either partially modify the text (Flux.1-Kontext-dev) or fail entirely. Additionally, models introduce unintended changes such as background color alterations and text sharpening effects. This demonstrates that text editing in historical documents remains an unresolved challenge for current models.
Historical document restoration [88] is a technique that aims to recover damaged or deteriorated historical document images, preventing further degradation and restoring their readability. Recently, some methods have explored using diffusion models [89, 90, 91] to accomplish this task. Common approaches require either pre-given missing content [89, 90] or predictions from specialized historical understanding models [91]. For simplicity, we do not provide the missing content and primarily test existing models’ repair ability of text and the damaged area. Results are shown in the bottom frame of Fig. 7. As observed, all models fail to produce text that resembles the original writing style and repair the damaged background texture. They mostly generate a new document rather than conducting “restoration”, demonstrating that existing generative models cannot solve such specialized OCR tasks so far.
3.2.3 Style Transfer and Super Resolution
Additionally, we evaluate the style transfer and super resolution tasks on historical documents, with results presented in Fig. 8. Regarding style transfer, all models fail to meet the target. GPT-4o reversed the target and source styles while generating incorrect content. Surprisingly, despite poor performance in previous tasks, Qwen-VLo-Preview is the only model that preserves the column structure of image 1 and the text sparsity of image 2, achieving the best relative output, although the text remains blurred and unreadable. Other models perform much worse than these two. For super resolution, GPT-4o and Qwen-VLo-Preview recognize the instruction of performing super resolution, but GPT-4o outputs a squared image with changed background color and Qwen-VLo-Preview distorts fonts. Other models seemingly fail to understand or misunderstand the prompt. For example, Flux.1-Kontext-dev actually blurs text rather than clarifying it. Super resolution is a common I2I translation task and we think that current generative models, especially those with trained understanding skills, should be able to perform this basic task. Yet, the results show the opposite and indicate significant room for improvement.
4 Result of Handwritten Text Image
Figure 9: Page-level T2I generation results for handwritten text images. Under the English scenario, most methods can generate accurate text of specific content (although some content may be missed). Qwen-VLo-Preview and Janus-4o are exceptions that generate unreadable and blurred text. However, under the Chinese scenario, only GPT-4o possesses the ability to generate Chinese text, while other models can not recognize the Chinese scripts and produce text in other languages or unreadable text. Flux.1-Kontext-dev even produces unrelated outcomes. For complete prompts, please refer to our GitHub repository.
Handwritten text images are ubiquitous in daily life, appearing in personal notes, letters, and exam papers. Unlike the uniformity of machine-printed fonts, handwriting exhibits significant diversity in style, slant, and hyphenation, reflecting the unique characteristics of individual writers. In recent years, with the rapid advancements in visual generation technologies, the task of handwritten text image generation has garnered substantial attention [63, 92, 93, 94, 95]. This technology not only enables personalized content creation and writing assistance but also provides a new avenue for augmenting training data for handwriting analysis tasks [96, 97, 98, 99, 100, 101]. Furthermore, the editing and removal of text in handwritten images hold broad practical applications in areas such as privacy protection, data cleaning, and educational scenarios [102, 103, 104]. Therefore, in this section, we evaluate the performance and potential of current generative models in three core tasks: handwritten text image generation, text editing, and text removal.
Figure 10: T2I generation results for handwritten text images in other content levels. Similar to the results in Fig. 9, while most models can successfully generate English handwritten text, they generally fail to generate Chinese handwritten lines or characters. Only GPT-4o and Qwen-VLO-Preview demonstrate some successful cases in paragraph or character generation. In addition, all models exhibit the capabilities of handling image-text interleaved scenarios, where GPT-4o performs best. For complete prompts, please refer to our GitHub repository.
4.1 T2I Generation
Generating handwritten text images from textual descriptions is a challenging text-to-image (T2I) generation task. This process demands that a model must accurately render the given textual content while simultaneously adhering to specific stylistic directives, including attributes like handwriting style (e.g., cursive, neat), ink color, and paper type. Consequently, this task serves as a robust benchmark for evaluating a model’s proficiency in both fine-grained text rendering and stylistic control. Our evaluation is structured across multiple levels of granularity: page-level, paragraph-level, line-level, character-level, and interleaved-level.
4.1.1 Page-Level
Fig. 9 presents the results for page-level handwritten text generation. Under the English scenario, most models can generate partially accurate text that is consistent with the prompt. GPT-4o stands out as the relatively best-performing model, producing clear, legible handwriting that closely matches specified styles such as “cursive” or “a slight right slant”. In contrast, Qwen-VLo-Preview can only render a small fraction of correct text, with the majority being unreadable. Janus-4o yields the poorest results, with its output being blurry and entirely illegible.
Figure 11: Text editing results for handwritten text images. The results show that editing specific handwritten text is a challenging task for existing generation models. They may repeat or lose some text that should be rendered (page-level, paragraph-level) and fail to maintain the original color and background (paragraph-level, line-level). A lot of room is still left for improvement.
The distinction in model capabilities becomes much more pronounced in the Chinese language scenario. Only GPT-4o demonstrates the ability to generate coherent and accurate Chinese text, successfully rendering complex characters in a handwritten style. The other models largely fail to recognize the Chinese script. They either produce unreadable scribbles or default to generating text in other languages. Notably, Flux.1-Kontext-dev exhibits a significant failure by generating a completely unrelated graphic instead of the requested text, indicating a fundamental misunderstanding of the prompt’s intent in a multilingual context.
4.1.2 Other Content Levels
We further evaluate the models on paragraph, line, character, and interleaved text-image generation, with the results presented in Fig. 10. At the paragraph level, most models exhibit significant difficulties with the Chinese paragraph prompt. GPT-4o is the only model to successfully generate a coherent and clear paragraph that accurately renders the text. As the task simplifies to the line and character levels, model performance improves, particularly in the English scenario. For the line generation task, GPT-4o again produces a nearly perfect and clear sentence, followed by OmniGen2. Other models only capture some keywords, while the rest of the text is poorly formatted. At the character level, most models can successfully generate the English letter "P". For the Chinese character, however, only GPT-4o and Qwen-VLo-Preview generate the correct character.
In the interleaved image-text scenario, which requires generating a diagram with explanatory text, GPT-4o demonstrates superior capabilities. It accurately generates a hand-drawn physics diagram of the law of reflection, complete with clear and correctly placed labels. Other models either produce inaccurate diagrams or render the textual labels illegibly.
4.2 Text Editing
We also evaluate the ability of models to edit handwritten text, a task that requires not only accurate understanding of editing instructions but also seamless integration of modifications while preserving the original handwriting style, background, and untouched content.
The performance of different models on text editing tasks is shown in Fig. 11. At the page level, although GPT-4o achieved relatively the best results, it did not perform a true "edit" but instead regenerated the entire page to incorporate the new text, leading to a loss of the original handwriting style and document texture. The performance of other models was even less satisfactory. For example, Qwen-VLo-Preview misinterpreted the prompt and inserted a stamp instead of text; Flux.1-Kontext-dev and OmniGen2 simply output the original text image without adding any new text; BAGEL introduced incorrect text; and Janus-4o produced completely chaotic and unusable results. At the paragraph and line levels, GPT-4o and Qwen-VLo-Preview successfully replaced the text, but failed to maintain the original writing style and untouched content, with issues such as content loss or errors, while other models failed entirely.
In summary, most current models lack the capability for fine-grained, context-aware editing. They often rely on a complete redrawing approach, which cannot retain the necessary style and background attributes of the original image. This highlights significant room for improvement in the task of handwritten text editing.
Figure 12: Results of the handwritten text removal task. Comparing the first and third rows, models excel at erasing text in a simple structure image like the case of the first row, while falling short in erasing handwritten text interleaved with other text (e.g., printed text). For removing specific handwriting like in the second and the fourth rows, models usually lack the ability to locate the target text and remove it with other text unchanged.
4.3 Handwritten Text Removal
We evaluate the handwritten text removal task, with results presented in Fig. 12. Two evaluation settings are adopted: erasing all text or erasing partial handwriting. Under the all-text removal case, given an image with simple structures (consisting of only one graphic, the text, and the background) as shown in the first row, most models can erase all handwritten text in the image. Conversely, in images with complex structures where handwritten text is embedded inside and interleaved with printed text, like the third row, they mostly fail to precisely remove the handwriting. On the other hand, as shown in the second and fourth rows, models generally fall short in removing the specific handwriting. They suffer from precisely locating the target text, mistakenly removing other text that should be preserved, and changing the structure or appearance of the background. This assessment suggests the shortage of current generative models in removing handwritten text, underscoring the necessity of future dedicated optimization.
5 Result of Scene Text Image
Figure 13: T2I generation results for scene text images. The red glowing indicates the relatively best output corresponding to each input. We test three language cases: English, Chinese, and mixed-language. Most models can handle English scene text generation. Yet, only GPT-4o and Qwen-VLo-preview handle Chinese text generation, Flux.1-Kontext-dev even delivers a totally unrelated result. For mixed-language, all models fail to handle this case, demonstrating chaotic, blurred, and nonsensical generated text. Even the relatively optimal-performing GPT-4o fails to generate all texts accurately, particularly small ones. For complete prompts, please refer to our GitHub repository.
Figure 14: Text editing and text removal results for scene text images. In scene text editing, while GPT-4o successfully modifies the target text, it adds unexpected text (the first row) and compromises the original aspect ratio. Other models exhibit inadequate text modification precision and substantial loss of textual and background texture fidelity. In scene text removal, models only fulfill the requirements in the cases of the first row, while accidentally eliminating other details like the animal and auxiliary text, or failing to erase smaller texts. This means scene text removal remains a challenge for current generative models.
Scene text, or text in the wild, refers to textual information that appears naturally in real-world environments. It can be found on product packaging, vehicle license plates, street signs, and numerous other objects in our daily surroundings. Unlike documents or handwriting, scene text is often characterized by its unconstrained font appearances and diverse backgrounds, such as curved text [105], multiple text orientations [106], perspective distortion [107], and low contrast and cluttered background [108]. Recent research has expanded beyond detection and recognition [96, 109, 110, 111] to focus on generative tasks [69, 62, 61, 112] involving scene text. These include T2I generation for synthesizing images with specific text content, scene text editing for modifying existing text while maintaining visual coherence, and scene text removal for erasing text while preserving background integrity. These generative applications support content creation, visual design, and privacy protection, though they remain challenging due to the complex integration of text within visual contexts. Therefore, we evaluate existing models on various scene text generation tasks to benchmark their capabilities in handling the complexity and diversity of text in real-world scenarios.
Figure 15: T2I generation results for images with naturally embedded text. As observed in the first and second rows, GPT-4o is pretty good at generating images with a small amount of text, perfectly following the instructions, while other models either fail to generate specified text or synthesize blurred, unclear text. However, in the third and fourth rows, all models can not accurately generate ordered text like numbers on the ruler or letters on the keyboard. This may be attributed to their lack of world knowledge or the insufficiency in synergizing their innate knowledge and generation performance. For complete prompts, please refer to our GitHub repository.
5.1 T2I Generation
The T2I generation result for scene text images is shown in Fig. 13. We present three cases featuring English, Chinese, and mixed-language text to comprehensively demonstrate the models’ scene text generation capabilities across different languages. Most models can accurately generate English scene text, with only BAGEL and Janus-4o failing in this case. However, only GPT-4o and Qwen-VLo-Preview can handle the Chinese scene text generation, while Flux.1-Kontext-dev produces a completely irrelevant image without any text, suggesting it may lack Chinese prompt understanding capability. For mixed languages, none of the models can accurately generate all the text, with most producing chaotic, blurred, or non-semantic content. We observe that although Qwen-VLo-Preview performs better in generating Chinese and English scene text, it struggles with other languages in the multilingual case, indicating potential limitations in this area, while GPT-4o demonstrates more balanced capabilities.
5.2 Scene Text Editing and Scene Text Removal
We also evaluate these models’ text editing and text removal capabilities on scene text images, as presented in Fig. 14. For text editing, only GPT-4o can accurately modify text according to instructions, but it also adds unexpected text (the first row) and fails to maintain the original aspect ratio of the input image. Other models fail to accurately modify the text and lose other textual or background texture details. For text removal, although some models can erase target text, many details remain problematic, such as changing background textures, failing to erase smaller text, or removing text and background that should be preserved. Moreover, we observed two main issues with these generative models when performing these two tasks: (1) Instruction understanding problems, where Flux.1-Kontext-dev cannot understand Chinese instructions and preserves the original image. (2) Text generation problems, where Janus-4o performs poorly in both modifying text and preserving the original text. In conclusion, scene text editing and scene text removal remain challenging for current models.
5.3 Naturally Embedded Text
Naturally embedded text refers to text that appears as an integral part of objects or scenes in images, where the text is seamlessly incorporated into the physical elements (e.g., smartwatch, keyboard) rather than being artificially overlaid. It is also a type of scene text, but differs from conventional scene text such as street signs and billboards, typically exhibiting greater diversity and complexity in terms of font, size, orientation, material, and integration methods. These texts often need to adapt to the functional requirements and physical constraints of the devices they appear on, such as display resolution, ergonomic keyboard design, or precise scale requirements of measuring tools, resulting in visual characteristics and recognition challenges distinctly different from ordinary scene text. Therefore, we specifically evaluate existing models’ T2I generation and text editing capabilities on naturally embedded text.
5.3.1 T2I Generation
We demonstrate the models’ T2I generation capabilities for naturally embedded text in Fig. 15. As observed in the first and second rows, GPT-4o performs excellently when generating electronic display fonts, and other models either fail to generate specified text or synthesize blurred, unclear text. However, when asked to generate items with ordered text, such as rulers or keyboards, all models struggle significantly; despite most being able to generate the correct items, they all exhibit problems with inaccurate ruler measurements and unreasonable keyboard characters. For instance, the rulers often show inconsistent or illogical measurement intervals and values, while the keyboards frequently display scrambled or repetitive characters instead of the standard QWERTY layout. This may be attributed to their lack of world knowledge or the insufficiency in synergizing their innate knowledge and generation performance, particularly in cases requiring precise spatial arrangement and sequential ordering of text elements.
Figure 16: Text editing results for images with naturally embedded text. Text editing in these cases involves fine-grained and small character modifications. From the results, GPT-4o performs best, successfully editing most texts to their target contents. However, it sometimes fails to preserve original text or pixel texture details (the second and third rows). Other models primarily struggle to modify small text (such as changing “5 km” to “63 km”) and sometimes output the vanilla input unchanged (OmniGen2 and BAGEL). Therefore, small text editing remains a significant challenge for existing generative models.
5.3.2 Text Editing
Finally, we evaluate the text editing task on naturally embedded text, with results shown in Fig. 16. We provide instructions for models to attempt to modify fine-grained and small text in images. Among all tested models, GPT-4o is the best-performing one, only failing when modifying Chinese text (the third row), though it lacks the ability to maintain the original text’s texture details. Other models primarily struggle to modify much smaller text (like changing "5 km" to "63 km") and sometimes directly output the vanilla input without changes (OmniGen2 and BAGEL). These results indicate that small text editing remains a significant challenge for existing generative models, particularly in cross-language application scenarios and maintaining visual consistency, providing clear directions for future optimization of multimodal models.
6 Result of Artistic Text Image
Artistic text is stylized typography that incorporates creative graphical components or decorative fonts to achieve aesthetic appeal beyond standard readable formatting. It has found widespread applications from creating memorable brand logos and advertisements to designing impactful movie titles, game interfaces, and personalized merchandise [113]. In the recent decade, numerous synthesis methods [114, 115, 116, 117] have been proposed to assist automatic glyph design, demonstrating remarkable font fidelity and diversity. Therefore, we incorporate this task into our evaluation, testing whether current generative models can fulfill this interesting, creative, and practical task.
6.1 T2I Generation
We first assess the T2I generation task, with results presented in Fig. 17. For line-level artistic text generation, the results are indeed inferior. Although existing models produce creative, visually pleasing artistic fonts, they suffer from content missing in both English and Chinese cases, which is not feasible for real-world applications. In addition, we impose on them to generate complex, rare Chinese characters, and they consistently fail to fulfill the requirements. The Flux.1-Kontext-dev stands as the worst performer, which generates human and photo frame images instead of text images. For character-level generation, all models can generate English single-character. However, for the rare Chinese character, they have failed again.
Figure 17: T2I generation results for artistic images. The outcomes of line-level text generation are inferior. All models, despite the relatively better GPT-4o, can not completely and accurately render long text in the artistic fonts. Also, they fail to generate images of rare Chinese characters. For complete prompts, please refer to our GitHub repository.
Figure 18: Text editing and style transfer results for artistic images. Overall, the editing and style transfer of the English script are well done by most models. Conversely, for the Chinese language, GPT-4o is the only model that can handle Chinese text editing and style transfer, which, however, still accidentally squares the image. This indicates that apart from the English language, multilingual text modification abilities need further improvement for current models.
6.2 Text Editing and Style Transfer
Subsequently, we assess the text editing and style transfer abilities of current generative models. Results are shown in Fig. 18. Here, style transfer looks similar to text editing due to their shared objectives of modifying the source text to the target. Hence, it is necessary to clarify the definitions of these two tasks: text editing requires the models to modify the text and keep other elements of the image unchanged, while style transfer aims at generating a new image with specific text content according to the source image’s text style, regardless of other elements in the vanilla images.
We evaluate text editing under two scenarios: including editing the text w/o and w/ style modifications. In the first and third rows, we instruct the model to modify text content while preserving the original style. All models demonstrate strong performance on English text, although GPT-4o, Qwen-VLo-Preview, and Janus-4o produce square outputs instead of maintaining the original rectangular aspect ratio. For Chinese text editing, only GPT-4o and Qwen-VLo-Preview show capability, but with notable artifacts: GPT-4o alters the background to black, while Qwen-VLo-Preview introduces unnecessary decorative elements. In the second row, most models successfully modifies both text content and writing style, while preserving background elements like fireworks and cityscape views. A minor issue is that some models turn the original rectangular images into squares.
As for style transfer, the source style are well transferred to the target content in the English cases. However, similar to text editing, most models can not handle style transfer of Chinese text. Collectively, these findings indicate that current generative models can handle English artistic text editing and style transfer effectively, but substantial improvements are needed for Chinese script processing.
7 Result of Complex and Layout-Rich Text Image
Complex and layout-rich text images, such as slides and posters, blend text, graphics, and intricate layouts to convey information effectively and aesthetically, which are widely used in marketing, education, and entertainment. The dynamic and context-sensitive nature of these layouts requires models to comprehend spatial arrangements, hierarchy, and visual balance. For instance, in a slide, text should align seamlessly with graphics to emphasize key points without causing clutter [118, 119, 120]. In addition, posters demand bold typography and striking visuals to capture attention [121, 122, 123, 124, 125]. Successfully generating and editing such images demands not only an understanding of textual semantics but also proficiency in design principles.
Figure 19: T2I generation and text editing results for slide images. GPT-4o demonstrates significantly better instruction-following abilities and generation quality in T2I generation than other models, which fulfills most requirements. However, other models fail to generate exact English/Chinese text but in unknown languages, also generating square images even given the instruction of generating a retangular one. Regarding text editing, actually all models fall short in this case, given rich and dense text like slide images. Locating fine-grained and a small portion of text to perform modification proves a great challenge for current generative models.
Figure 20: T2I generation and text editing results for poster images. Most models can handle the generation of English posters, whereas they fall short in generating Chinese posters, primarily attributed to their lack of Chinese training data. Similar to Fig. 19, current models mostly cannot accurately modify text, and even when they can modify text, they cannot guarantee that other original details remain unchanged.
Figure 21: Results of the layout-aware text generation task. For English cases, OmniGen2 is the only one that adds correct text in an appropriate position without obstructing the original main component and preserving other pixel details. For Chinese cases, GPT-4o stands as the only model that adds correct text in appropriate positions. Other models seem to lack the ability to comprehend and generate Chinese text.
7.1 Slide Image
7.1.1 T2I Generation and Text Editing
Typically, slide images consist of rich and structured elements such as headings, text blocks, and graphics, ordered in a hierarchical layout. This presents great challenges to generation models for comprehending both textual semantics and design aesthetics in slide creation. The generated results are presented in the top frame of Fig. 19. As observed, while GPT-4o suffers from text errors, it demonstrates significantly better instruction-following abilities than other models, producing readable text and more appealing layouts. In contrast, other models generate illegible text and sometimes output totally unrelated images. For example, Qwen-VLo-Preview produces an image of handwritten text but a slide. Similarly, Flux.1-Kontext-dev fails to interpret Chinese instructions, instead generating an unrelated image of a young girl. These results highlight the challenges generative models face in producing complex slide content.
We also evaluate their text editing capabilities of slide images, with results illustrated in the bottom part of Fig. 19. It is observed that all models showcase unsatisfactory performance. GPT-4o partially executes the requested edits but alters portions of the original layout and modifies accompanying images. Flux.1-Kontext-dev similarly fails to locate the source text and preserve the layout. Qwen-VLo-Preview and OmniGen2 introduce extensive changes to background colors and text, producing unreadable text. BAGEL fails to respond to the instructions and leaves the image unchanged, while Janus-4o even produces chaotic content with no sense. Given rich and dense text in the slide, models struggle to locate a small portion of text while also preserving other elements.
7.2 Poster Image
7.2.1 T2I Generation and Text Editing
Similar to slide images, poster stands as an important type of visually rich image that deserves investigation. The results of poster generation are shown in the top frame of Fig. 20. GPT-4o delivers the most consistent results, correctly and clearly rendering all English text, despite failing to render some small Chinese text as shown in the third row. Other models demonstrate certain abilities in text generation, while the results are often erroneous or incomplete, especially for Chinese text. Flux.1-Kontext-dev even produces a gate image instead of a poster. Models like OmniGen2, BAGEL, and Janus-4o frequently produce unreadable or nonsensical text, rendering their outputs unsuitable for practical use.
In text editing, as illustrated at the bottom of Fig. 20, GPT-4o manages to correctly modify partial requested content but makes unintended changes to other parts of the poster, such as altering design elements or unrelated sections of text. Qwen-VLo-Preview partially handles Chinese text editing but often generates incomplete or incorrect modifications. The remaining models either fail to edit Chinese text altogether or produce outputs with unreadable results. None of the models, apart from GPT-4o, successfully locate and edit the correct target text in English, and all exhibit issues with altering other design elements during the editing process.
7.3 Layout-aware Text Generation
Layout-aware text generation, also known as content-aware layout generation [126, 127, 128, 129, 130], requires models to position texts naturally within images, ensuring they do not oblige the original graphical component for visually appealing. The evaluation results are shown in Fig. 21. GPT-4o is the best performer that correctly adds text without compromising the original layouts in the last three rows, while it inadvertently alters the image in the first row. Qwen-VLo-Preview successfully perceives the image layout for appropriate text placement but delivers incomplete or unreadable text rendering. Other models primarily produce erroneous or readable text, and sometimes leave the images unchanged. Specifically, Flux.1-Kontext-dev suffers from illumination of generating two lines of “Camera is good”, and OmniGen2 even generates English text given a Chinese prompt (the last row). The results prove the challenge of this task for current generative models.
8 Conclusion
In this paper, we comprehensively evaluate the image generation/editing capabilities of several SOTA generative models in a spectrum of OCR generative tasks. 33 common OCR tasks are chosen for evaluation, which are then categorized into five main categories based on text characteristics. We subsequently select six cutting-edge generative models, including two closed-source models, i.e., GPT-4o [57] and Qwen-VLo-Preview [60], and four open-source models, i.e., Flux.1-Kontext-dev [30] and OmniGen2 [31] (specialized generation models), as well as BAGEL [50] and Janus-4o [44] (unified understanding and generation models). Our evaluation reveals that text image generation and editing remain significant challenges for current up-to-date models, in which they suffer from inaccurate text locating, poor structural preservation, blurred or illegible character generation, etc. Given the lack of dedicated evaluation for image generation performance in the OCR field, our work fills this gap for the first time and provides crucial insights into current generative models’ limitations and potential future directions. We hope our analysis can provide useful insights for the community and inspire increased efforts to improve text image synthesis and editing, internalizing high-quality text image generation skills into general-domain generative models to advance our step toward AGI.