### v3

# Aesthetics is Cheap, Show me the Text: An Empirical Evaluation of State-of-the-Art Generative Models for OCR

Abstract.
Text image is a unique and crucial information medium that integrates visual aesthetics and linguistic semantics in modern e-society. Due to their subtlety and complexity, the generation of text images represents a challenging and evolving frontier in the image generation field. The recent surge of specialized image generators (e.g., Flux-series) and unified generative models (e.g., GPT-4o), which demonstrate exceptional fidelity, raises a natural question: can they master the intricacies of text image generation and editing? Motivated by this, we assess current state-of-the-art generative models’ capabilities in terms of text image generation and editing. We incorporate various typical optical character recognition (OCR) tasks into our evaluation and broaden the concept of text-based generation tasks into OCR generative tasks. We select 33 representative tasks and categorize them into five categories: document, handwritten text, scene text, artistic text, and complex & layout-rich text. For comprehensive evaluation, we examine seven models across both closed-source and open-source domains, using tailored, high-quality image inputs and prompts. Through this evaluation, we draw crucial observations and identify the weaknesses of current generative models for OCR tasks. We argue that photorealistic text image generation and editing should be internalized as foundational skills into general-domain generative models, rather than being delegated to specialized solutions, and we hope this empirical analysis can provide valuable insights for the community to achieve this goal. This evaluation is online at https://github.com/NiceRingNode/Awesome-Generative-Models-for-OCR.
OCR, Generative Models, Unified Understanding and Generation
$c$s: Human-centered computing Empirical studies in visualization
1. Introduction
images with machines represents humanity’s ambitious pursuit to translate the visual world into algorithms, teaching computers the fundamental human skill of creation. Since the 1960s, researchers have started to use computers for creating artworks like films and still ASCII art (Noll, 2016; Krampen and Seitz, 1967; Goodchild, 2021; Computer History Museum, 2024). Later, this field evolved from early human-programmed, procedural rules (Cohen, 1973; Perlin, 1985) to modern, data-driven artificial intelligence (AI) approaches, exemplified by the early Variational AutoEncoders (VAE) (Kingma et al., 2013; Higgins et al., 2017) and Generative Adversarial Networks (GAN) (Goodfellow et al., 2014; Karras et al., 2019). Beyond image-to-image translation or conditional image generation (Elasri et al., 2022), researchers sought to control the generation process more naturally and flexibly. This fueled the emergence of text-to-image (T2I) generation (Reed et al., 2016), allowing users, even those non-technical, to generate visual content with natural language. Building upon the proliferation of T2I generation (Qiao et al., 2019; Li et al., 2019; Nonghai Zhang, 2024), the community has now broaden the application realm into instruction-guided image generation and editing, driven by the confluence of large-scale text-image datasets, advances of language models (Radford et al., 2021; Raffel et al., 2020), and innovative architectures like diffusion (Ho et al., 2020; Rombach et al., 2022), flow matching (Lipman et al., 2023), and autoregressive generation (Chen et al., 2020; Li et al., 2024a).
Mainstream state-of-the-art (SOTA) generation approaches fall into two types: specialized image generation models and native unified generative models. The first type is primarily built based on diffusion or flow matching framework, such as DALL·E (Ramesh et al., 2021, 2022; OpenAI, 2023), Imagen (Saharia et al., 2022), Stable Diffsuion (Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024), Flux (Labs, 2025; Labs et al., 2025), and OminGen (Wu et al., 2025d; Xiao et al., 2025). They focus merely on generating or editing images based on instructions, using a relatively small text encoder (Radford et al., 2021) for textual prompt encoding. In contrast, the second type unifies image understanding and generation within one Transformer model, termed unified generative models. They are typically built based on multimodal large language models (MLLMs) with the autoregressive objective, which can be further divided into three meta-categories. (1) Models w/o diffusion modeling. This line of models uses VQGAN (Esser et al., 2021) or VQVAE (van den Oord et al., 2017) to tokenize images into discrete tokens as input to the language models for multimodal autoregressive learning, and detokenizes the output features back into images (Lu et al., 2024; Team, 2024; Liu et al., 2025; Wu et al., 2025a; Ma et al., 2025; Wu et al., 2024; Wang et al., 2024; Chen et al., 2025b). (2) Diffusion-embedded models. In contrast, this type of works embrace diffusion modeling, integrating diffusion’s loss and the iterative denosing process into the Transformer for training and inference, respectively (Zhou et al., 2024; Xie et al., 2025a, b; Zhao et al., 2024; Shi et al., 2024; Deng et al., 2025). They can switch their role as language models for autoregressive text generation or diffusion models for parallel image synthesis. (3) Models with extra diffusion head. Unlike the second meta-category, this fashion of approach retains only autoregressive next-token prediction loss for training while incorporating an additional, lightweight diffusion head (Li et al., 2024a) alongside the MLLM, which is conditioned on the MLLM’s hidden states for image generation (Fan et al., 2025; Tong et al., 2024; Sun et al., 2024; Pan et al., 2025; Wu et al., 2025c). Recently, the release of Gemini-2.0-Flash (DeepMind, 2025) and GPT-4o (OpenAI, 2025) marks a milestone of unified native image generation, exhibiting commercial-level image generation and editing capabilities without compromising understanding skills.
Figure 1. Task categorization and subordination. We classify 33 OCR generative tasks into five primary categories based on text characteristics: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text. Each primary category encompasses multiple sub-tasks.
Text images, such as documents, scene texts, and handwriting, represent special and critical visual information carriers for human communication. Unlike natural objects or paintings, text images contain structured linguistic information that requires both visual perception and language understanding for proper interpretation. This dual requirement makes the generation and manipulation of text images a particularly challenging frontier in the generative AI community (Elanwar and Betke, 2025; Ren et al., 2025). Despite significant advances in large-scale T2I synthesis, generating high-quality text images remains problematic due to fundamental challenges, such as ensuring precise character formation and legibility across scales, maintaining semantic consistency between text and visual content, and handling diverse layouts and orientations. General-domain generators excel at natural image generation but usually falter in text image generation and editing, and very few of them (Esser et al., 2024; OpenAI, 2025; Team, 2025) have formally acknowledged or addressed this critical deficiency. We argue that photorealistic text generation is an important and indispensable ability for achieving artificial general intelligence (AGI). Although numerous models tailored for text image synthesis have been proposed (Chen et al., 2023; Tuo et al., 2024; Zhu et al., 2023; Luo et al., 2023; Yang et al., 2024; Wang et al., 2025), text generation should be internalized as a foundational skill, rather than delegated to specialized solutions. In addition, despite numerous evaluation on diverse content generation fields, such as face/code/human motion generation, (Kammoun et al., 2022; Schäfer et al., 2024; Zhu et al., 2024a; Fakhoury et al., 2024), a comprehensive assessment dedicated to text image generation/editing is lacking. Such an assessment is essential for systematically evaluating the strengths and weaknesses of existing models, illuminating clear pathways for future research.
To this end, we conduct a comprehensive evaluation of text image generation and editing capabilities in SOTA generative models. In literature, the automatic interpretation of text images is known as optical character recognition (OCR) (Islam et al., 2017), among which many tasks are inherently generative, such as document deshadowing (Zhang et al., 2024b), scene text removal (Liu et al., 2020b), and font generation (Yang et al., 2024). Hence, we broaden the scope of text generation and reframe these problems as OCR generative tasks. Our evaluation operates in three stages. First, we curate 33 OCR tasks and classify them into five main categories based on text properties: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text. Second, we select several representative models for benchmarking, including both specialized generators and unified models from closed-source and open-source domains: GPT-4o (OpenAI, 2025), Qwen-VLo-Preview (Team, 2025), Qwen-Image (Wu et al., 2025b), Flux.1-Kontext-dev (Labs et al., 2025), OmniGen2 (Wu et al., 2025d), BAGEL (Deng et al., 2025), and Janus-4o (Chen et al., 2025b). Third, we tailor expressive prompts and select high-quality input images from public datasets (Chng et al., 2019; Hradiš et al., 2015) or the web for each task, encompassing bilingual evaluation in both English and Chinese. From the evaluation, we discover that text image generation and editing remain substantial challenges for existing generative models, and draw several crucial observations as follows.
  1. (1)
    Awesome creativity. Existing models excel at generating creative and design-oriented images with text, especially posters, slides, and street scenes, when given detailed prompts. Regardless of the text quality, the style, color harmony, and overall composition often demonstrate remarkable artistic appeal and sophistication.
  2. (2)
    Insufficient perception and localization for specific text. In text removal and editing tasks, current models struggle to accurately modify designated regions, often resulting in incomplete modifications or unintended alterations in other areas. This prohibits them from performing fine-grained text content manipulation.
  3. (3)
    Poor structural preservation. While some tasks require modifications to specific text regions, current models often fail to preserve other areas that should remain unchanged. For example, in text removal, editing, and layout-aware text generation, models may successfully manipulate the target text but inadvertently alter the style, position, or content of surrounding text and background details. Similarly, document enhancement operations such as dewarping, deshadowing, and deblurring frequently introduce unwanted changes to document layout and content beyond the intended corrections.
  4. (4)
    Unstable instruction-following ability and hallucination. Models occasionally fail to follow user instructions and produce unstable outputs. Flux.1-Kontext-dev frequently generates entirely unrelated images in T2I tasks. Qwen-VLo-Preview (document dewarping), OmniGen2 (document dewarping, document deblurring, and modern/historical document editing), and BAGEL (appearance enhancement) occasionally produce anomalous outputs that deviate from user instructions.
  5. (5)
    Problematic processing of complex content. Existing models typically struggle with complex visual content. In document dewarping, embedded graphics cannot be properly restored. In document deshadowing, tables with hierarchical text and sophisticated structures can be accurately replicated. Also, while they can generate simplified Chinese characters, they cannot handle complex Chinese characters.
  6. (6)
    Suboptimal synthesis of dense, long text images. As the required amount and density of generated text or the text inside the input image increase, the results become more prone to errors. In document-related tasks, the generated text is sometimes blurred due to high density. Paragraph-level handwriting generation showcases more errors compared to line-level results. This phenomenon is further exacerbated given complex fonts, as exemplified in generating long text with artistic font styles.
  7. (7)
    Constrained resolution and aspect ratio control. Outputting correctly edited images while maintaining the original resolution or aspect ratio remains challenging. While GPT-4o excels at image editing (except handwritten images), it outputs dimensions in 512-pixel multiples, inevitably distorting image layouts. Other models like BAGEL and OmniGen2 preserve image aspect ratios but demonstrate weaker editing capabilities.
  8. (8)
    Deficient unified understanding and generation skill. Unified generative models, by virtue of training on visual understanding tasks, are expected to develop better comprehension skills. Yet, except for GPT-4o, the other three models exhibit limited world knowledge to fully and precisely interpret OCR-related instructions, particularly evident in professional tasks like dewarping, deblurring, and historical document restoration. Larger scale of training data for OCR-based analysis and text-centric feature input may alleviate this issue.
  9. (9)
    To-be-improved vertical-domain task performance. For highly specialized vertical-domain tasks, such as modern document processing tasks and historical document restoration, these general-domain models demonstrate inferior proficiency. Incorporating these tasks into model training could improve both general and specialized generation capabilities.
  10. (10)
    Limited multilingual proficiency. While current models handle English text generation, many fail to generate Chinese text, particularly open-source ones. The generated Chinese text is often garbled, illegible, or missing characters, and instructions specifying Chinese text are sometimes incorrectly rendered in other languages. Adding multilingual support is essential for universally and linguistically applicable generative models.
  11. (11)
    Large gap between open-source and closed-source models. In terms of text generation and editing, closed-source models, especially GPT-4o, significantly outperform the open-source ones. While this could be partially limited by the smaller size of open-source models (7B), more future efforts should be devoted to improving their text synthesis skills. Scaling laws specifically for text image generation have not been explicitly unveiled, representing a promising research direction.
Notably, while some prior works (Chen et al., 2025a; Cao et al., 2025; Li et al., 2025) have evaluated closed-source unified generative models’ (particularly GPT-4o) generation abilities in general visual domains, we transition the focus to text images and expand our evaluation to both open-source and closed-source models. We hope this work can inspire broader research efforts toward improving visual text generation and editing, establishing these skills as fundamental capabilities rather than specialized niches.
2. Evaluation Setting
Evaluated tasks. We curate 33 common OCR tasks that can be formulated as generation or editing objectives, including document dewarping, handwritten text removal, scene text editing etc. These tasks are categorized into five main categories: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text, where each main category contains multiple sub-tasks. For each sub-task, we design corresponding English and Chinese prompts as consistent inputs across all tested models. A mindmap introducing detailed task subordination is shown in Fig. 1.
Tested models. Our evaluation considers both closed-source models and open-source models. Closed-source models include GPT-4o (OpenAI, 2025) and Qwen-VLo-Preview (Team, 2025); while open-source models include Qwen-Image (Wu et al., 2025b), Flux.1-Kontext-dev (Labs et al., 2025), OmniGen2 (Wu et al., 2025d), BAGEL (Deng et al., 2025), and Janus-4o (Chen et al., 2025b). Except for Flux.1-Kontext-dev and OmniGen2, which are specialized image generation models, all others are unified generative models that are capable of both understanding and generation. In addition, all models support text-to-image (T2I) generation and instruction-based image editing, competent for our evaluation tasks. An overview of these models is summarized in Table 1.
Evaluation. We present qualitative results through visualization for each task. Quantitative metric computation is temporarily not considered. Each visualization presents the used prompts, input images (if any), and output images from all models in a grid format for intuitive comparisons.
Table 1. Overview of tested models. Pub. denotes publication. Und. denotes understanding. Gen. denotes Generation. GPU Memory specifically refers to GPU memory requirements for model inference.
Model Pub. Date Type Availability #Parameters Max Context Length GPU Memory
GPT-4o (OpenAI, 2025) 2025.3.25 Unified Und. & Gen. Closed-Source - - -
Qwen-VLo-Preview (Team, 2025) 2025.6.26 Unified Und. & Gen. Closed-Source - - -
Qwen-Image (Wu et al., 2025b) 2025.8.18 Specialized Gen. Open-Source 20B 32,768 tokens 3 $\times$ 48GB
Flux.1-Kontext-dev (Labs et al., 2025) 2025.6.17 Specialized Gen. Open-Source 12B 77 tokens 48GB
OmniGen2 (Wu et al., 2025d) 2025.6.23 Specialized Gen. Open-Source 7B 128,000 tokens 48GB
BAGEL (Deng et al., 2025) 2025.5.20 Unified Und. & Gen. Open-Source 7B 32,768 tokens 2 $\times$ 48GB
Janus-4o (Chen et al., 2025b) 2025.6.22 Unified Und. & Gen. Open-Source 7B 4,096 tokens 48GB
3. Result of Document Image
3.1. Modern Document
Figure 2. Dewarping results for modern documents. The red glowing indicates the relatively best output corresponding to each input. The overall dewarping results are inferior, where GPT-4o rectifies the image to be flat but loses some embedded graphics and text, while other methods even fail to perform dewarping and lose substantial textual details.
Modern documents play an essential role in daily human activities for information transmission, consisting of text, graphics, tables, and other informational elements. Hence, document digitization and processing have become increasingly crucial for efficient information management and automated analysis, with OCR serving as a core technique (Tang et al., 1996; Subramani et al., 2020). Typically, the digitized documents require preprocessing before OCR to enhance accuracy, including flattening the scrumpling of the documents, removing shadows, and sharpening the text. These preprocessing technologies constitute different document processing tasks, most of which can be formulated as an image-to-image translation manner. Given the recent trend of instruction-based image generation and editing, we investigate whether these models can address important and practical document-related tasks.
Figure 3. Deshadowing and deblurring results for modern documents. Flux.1-Kontext-dev exhibits the optimal deshadowing results in both language cases, with most texts preserved and shadow removed. Other models either fail to remove the shadow or mistakenly repeat the textual content. For document deblurring, Flux.1-Kontext-dev showcases nearly perfect results in the English scenario. GPT-4o performs better in the Chinese case with precise text restoration but fails to restore the document structure.
Figure 4. Appearance enhancement results for modern documents. Only GPT-4o and Flux.1-Kontext-dev successfully interpret the instruction of outputting “PDF-like” documents. Most models fail to preserve document structures, particularly evident in erroneous repetition of table contents.
3.1.1. Document Dewarping
Document dewarping refers to correcting geometric distortions of document images, typically caused by curved, warped, or folded pages during scanning or photographing. Models are required to output a flattened document surface with preserved original text, facilitating reading or further OCR processing (Zhang et al., 2022b; Feng et al., 2025; Zhang et al., 2025b, 2026). Existing dewarping methods compute displacement fields that map the warped input image to the dewarped outputs. Yet, from the image generation perspective, this can be formulated as an image-to-image translation task. Early generative models could not interpret the “dewarping” or “rectification” instructions nor handle the drastic, detailed pixel changes to derive an accurate, flattened document. With advances in image generators, especially unified generative models capable of both understanding and generation, they may now understand the “dewarping” instruction and conduct sophisticated pixel transformation. More broadly, this task somehow represents the perception and comprehension abilities of real-world physics, motivating our detailed evaluation.
Results of the document dewarping task are presented in Fig. 2. All the dewarped results prove unsatisfactory. GPT-4o is the only model that identifies the “dewarping” instruction and produces flattened output documents, showcasing impressive instruction-following capability. While this is surprising, GPT-4o loses substantial portions of original text content and some embedded graphics, and also fails to render Chinese content, making the dewarped output infeasible for real-world usage. Other models mostly fail to understand and follow the “dewarping” or “crop” instructions, generating blurred, chaotic text. Qwen-VLo-Preview even exhibits hallucination without following the input and instructions, as shown in the penultimate row, third column of Fig. 2.
Figure 5. Text editing results for modern documents. GPT-4o and Qwen-VLo-preview better follow the instructions compared to other unified models, especially in locating target text and conducting replacement. Flux.1-Kontext-dev significantly outperforms other specialized generation models.
3.1.2. Document Deshadowing
Document deshadowing refers to removing shadows from document images (Li et al., 2023a). Similar to document dewarping, we require models to output deshadowed or deblurred images according to the input image and instructions. Results are presented in the top panel of Fig. 3. For deshadowing, Flux.1-Kontext-dev demonstrates optimal performance by effectively removing shadows while preserving original text content and background color. Other models either fail to remove the shadow (Qwen-Image, OmniGen2, BAGEL), inadvertently repeat text content, or alter the document’s background color (GPT-4o, Qwen-VLo-Preview). Qwen-Image even delivers a mosaic-like output rather than a document as shown in the first row.
3.2. Document Deblurring
Document deblurring refers to removing blur artifacts from document images to restore sharp, readable text and clear visual content (Souibgui and Kessentini, 2022). Results are presented in the bottom panel of Fig. 3. As observed, English text proves easier to process than Chinese text. Flux.1-Kontext-dev delivers nearly perfect results with clear text and correct content positioning, while GPT-4o seems to automatically complete content and generate a new image, possibly witnessing this document’s content during training. OmniGen2 may misunderstand the instruction as erasing all text, producing blank images. With Chinese text, all models fail to deblur the text while keeping original content structure (title, subtitle, main text). Surprisingly, although not a unified understanding and generation model, Flux.1-Kontext-dev demonstrates notable low-level visual document processing abilities, showcasing potential for further development. In contrast, Qwen-Image outputs drawing-like or unreadable documents, suggesting its shortcomings in processing documents with dense text.
3.2.1. Document Appearance Enhancement
Document appearance enhancement, also known as document illumination rectification (Feng et al., 2021; Zhang et al., 2024a), refers to the technique of mainly correcting uneven lighting conditions while also removing other real-world degradations like shadows and bleed-through to improve camera-captured documents’ visual quality. Evaluation results are presented in Fig. 4. In the first row, while Flux.1-Kontext-dev achieves relatively good enhancement and text preservation, it adds unexpected woodgrain edges. BAGEL fails to correct the lighting and outputs a red-green-mixed background. Qwen-VLo-Preview, Qwen-Image, and Janus-4o generate chaotic, unreadable text. GPT-4o and OmniGen2 exhibit erroneous text repetition similar to the document deshadowing task. In the second row, no model can perfectly preserve the table content, and the restoration of Chinese text is unsatisfactory. Furthermore, only GPT-4o and Flux.1-Kontext-dev comprehend the instruction to output “PDF-like” documents, suggesting the need to improve generative models’ understanding of professional document terminology.
3.2.2. Text Editing
We also evaluate text editing capabilities on PDF or real-world documents, with results shown in Fig. 5. GPT-4o, Qwen-VLo-Preview, and Flux.1-Kontext-dev can follow instructions to perform target modifications, although the modifications are sometimes incomplete or incorrect. However, other models consistently fail to edit documents. OmniGen2 and BAGEL fail to locate the target text and perform no/wrong alterations. Even worse, Qwen-Image and Janus-4o can not preserve the original text content but produce chaotic, unreadable content. As specialized generation models, Flux.1-Kontext-dev significantly outperforms Qwen-Image despite a smaller model scale (13B vs 20B), making it a more feasible option for text-editing fine-tuning in document scenarios due to superior performance and lower training costs.
3.3. Historical Document
Figure 6. T2I generation results for historical documents. GPT-4o fulfills generation requirements of specific page number and the synthesis of both English and Chinese text, despite sometimes the content is incomplete or incorrect. However, other models can generate ancient scripts but mostly fail to produce readable textual content. For complete prompts, please refer to our GitHub repository.
Historical documents represent invaluable repositories of human cultural heritage. The digitization and analysis of historical documents (Philips and Tabrizi, 2020; Nikolaidou et al., 2022; Cao et al., 2024) has emerged as a critical research area, aiming to aid in the preservation and understanding of these ancient cultural artifacts. These documents contain dense textual content, complex layouts, and antiquated fonts, which present a great challenge to image generation models. Therefore, we benchmark existing models on various historical document-related tasks, including general tasks like T2I generation and text editing, as well as vertical-domain tasks like historical document restoration.
Figure 7. Text editing and document restoration results for historical documents. For text editing, all models fail to modify specific Chinese text. The relatively best GPT-4o suffers from repeated and wrong editing, while the worst Qwen-Image even generates new images with specified text content, deviating from the editing objective. Yet, some models like GPT-4o and Flux.1-Kontext-dev demonstrate promising edit results on English text editing. For historical document restoration, all models fail to generate new text while preserving the original writing style and maintaining consistent backgrounds.
3.3.1. T2I Generation
We present T2I generation results for historical documents in Fig. 6. We design detailed prompts describing the text content, font style, writing paper details, and other environmental factors such as desks and oil lamps. We ask models to generate one-page historical scripts in the first three rows and a three-page script in the fourth row. For instruction following, most models successfully generate ancient script, albeit sometimes not one-page or three-page. Conversely, Flux.1-Kontext-dev produces irrelevant outcomes like a book or human figures, suggesting it may lack historical document training data. Regarding text rendering, while models can somehow generate English text, they consistently fail to render Chinese text. GPT-4o is the only model capable of generating both English and Chinese fonts while following page requirements correctly. Nevertheless, GPT-4o sometimes produces incomplete or incorrect text.
Figure 8. Style transfer and super resolution results for historical documents. For style transfer, while Qwen-VL-Preview produces the relatively best result by preserving the column structure from image 1 and text sparsity from image 2, the rendered text is blurred and lacks semantics. Other models can not follow the instructions and produce unreadable results. In the super resolution task, GPT-4o and Qwen-VLo-Preview improve image resolution but at the cost of color change or font distortion, while the others misunderstand the instructions (Flux.1-Kontext-dev) or generate completely unrelated results (Qwen-Image).
3.3.2. Text Editing
We then perform text editing on historical documents, with results shown in the top frame of Fig. 7. For Chinese text editing, tested models consistently fail to modify Chinese text images. Although GPT-4o locates and attempts to modify the target text, the edited content is incorrect and the nearby text is accidentally altered. For English text, GPT-4o and Qwen-VLo-Preview successfully render the target text “COVERING”, but only GPT-4o places it in the correct place. Other models either partially modify the text (Flux.1-Kontext-dev) or fail entirely. Notably, Qwen-Image generates brand-new images rather than editing the original image, indicating its poor editing ability under this scenario. In addition, models introduce unintended changes such as background color alterations and text sharpening effects. This demonstrates that text editing in historical documents remains an unresolved challenge for current models.
3.3.3. Historical Document Restoration
Historical document restoration (Hedjam and Cheriet, 2013) is a technique that aims to recover damaged or deteriorated historical document images, preventing further degradation and restoring readability. Recently, some methods have explored using diffusion models (Zhu et al., 2024b; Yang et al., 2025; Zhang et al., 2025c) for this task. Common approaches require either pre-given missing content (Zhu et al., 2024b; Yang et al., 2025) or predictions from specialized historical understanding models (Zhang et al., 2025c). Here, we do not provide missing content and primarily test existing models’ repair abilities of text and damaged areas. Results are shown in the bottom frame of Fig. 7. All models fail to preserve original writing style or repair damaged background texture. They mostly generate new documents rather than conducting “restoration”, indicating that existing generative models cannot solve such specialized OCR tasks and demand dedicated optimization.
3.3.4. Style Transfer and Super Resolution
Additionally, we evaluate the style transfer and super resolution tasks on historical documents, with results presented in Fig. 8. Regarding style transfer, all models fail to meet the target. GPT-4o reversed the target and source styles while generating incorrect content. Surprisingly, despite poor performance in previous tasks, Qwen-VLo-Preview is the only model that preserves the column structure of image 1 and the text sparsity of image 2, achieving the best relative output, although the text remains blurred and unreadable. Other models perform much worse than these two.
For super resolution, GPT-4o and Qwen-VLo-Preview recognize the instruction of performing super resolution, but GPT-4o outputs a squared image with a changed background color and Qwen-VLo-Preview distorts fonts. Other models misunderstand or entirely fail to comprehend the prompt. Flux.1-Kontext-dev actually blurs text instead of enhancing clarity, while Qwen-Image generates an unrelated map image rather than improving resolution, further demonstrating its limitations in image editing tasks. Super resolution is a common image-to-image translation task and we think that current generative models, especially those with trained understanding skills, should be able to perform this basic task. Yet, the results reveal significant deficiencies and highlight substantial room for improvement in existing models.
4. Result of Handwritten Text Image
Figure 9. Page-level T2I generation results for handwritten text images. For English, most methods can generate accurate texts of specific content (although some content may be missed). Qwen-VLo-Preview and Janus-4o are exceptions that generate unreadable and blurred text. However, for Chinese, GPT-4o and Qwen-Image exhibit the ability to generate Chinese text, while other models can not recognize Chinese scripts and produce text in other languages or unreadable text. Flux.1-Kontext-dev even produces unrelated outcomes. For complete prompts, please refer to our GitHub repository.
Handwritten text images are ubiquitous in personal notes, letters, and exam papers. Unlike the uniformity of machine-printed fonts, handwriting exhibits significant diversity in style, slant, and hyphenation, reflecting the unique characteristics of individual writers. Recent advancements in visual generation technologies have brought increasing attention to handwritten text image generation (Zhu et al., 2023; Das et al., 2023; Dai et al., 2024; Pippi et al., 2025). This technology enables writing assistance (Ha and Eck, 2018; Schäfer et al., 2023) and augments training data for handwriting analysis tasks (Luo et al., 2019; Zhang et al., 2022a; Peng et al., 2022; Yang et al., 2023; Zhang and Jin, 2024; Zhang et al., 2025a). Additionally, the editing and removal of text in handwritten images hold broad practical applications in areas such as privacy protection, data cleaning, and educational scenarios (Li et al., 2024b; Huang et al., 2023; Wang et al., 2022). We therefore evaluate the performance of current models in three tasks: handwritten text image generation, text editing, and text removal.
Figure 10. T2I generation results for handwritten text images in other content levels. Similar to Fig. 9, while most models successfully generate English handwritten text, they generally fail with Chinese handwritten lines or characters. Only GPT-4o and Qwen-VLo-Preview demonstrate some successful cases in paragraph or character generation. In addition, all models exhibit capabilities for handling image-text interleaved scenarios, with GPT-4o performing best. For complete prompts, please refer to our GitHub repository.
Figure 11. Text editing results for handwritten text images. The results show that editing specific handwritten text is a challenging task for existing generation models. They may repeat or lose some text that should be rendered (page-level, paragraph-level) and fail to maintain the original color and background (paragraph-level, line-level). A lot of room is still left for improvement.
Figure 12. Text removal results for handwritten text images. Comparing the first and third rows, models excel at erasing text in simple structure images, while falling short in erasing handwritten text interleaved with other text (e.g., printed text). For removing specific handwriting, as shown in the second and fourth rows, models typically fail to accurately locate target text while preserving surrounding content.
4.1. T2I Generation
Generating handwriting images with textual descriptions is a challenging T2I generation task. It demands models to accurately render given textual content while adhering to specific stylistic directives, such as writing style, ink color, and paper type. We evaluate this task across multiple granularity levels: page-level, paragraph-level, line-level, character-level, and interleaved-level.
4.1.1. Page-Level
Fig. 9 presents the results for page-level handwritten text generation. Under the English scenario, most models can generate partially accurate text consistent with the prompt. GPT-4o performs best, producing clear, legible handwriting that closely matches specified styles such as “cursive” or a “slight right slant”. Qwen-Image also performs well by producing accurate text, albeit it does not follow the specified writing style instructions. In contrast, Qwen-VLo-Preview can only render a small fraction of correct text, with the majority being unreadable. Janus-4o yields the poorest results, with its output being blurry and entirely illegible.
The distinction in model capabilities becomes much more pronounced in the Chinese language scenario. GPT-4o and Qwen-Image demonstrate the ability to generate coherent and accurate Chinese handwritten text. The other models largely fail to recognize the Chinese script. They either produce totally unreadable scribbles or default to generating text in other languages. Notably, Flux.1-Kontext-dev exhibits a significant failure by generating a completely irrelevant graphic instead of the requested text, indicating the inability to perceive the prompt’s intent.
4.1.2. Other Content Levels
We further evaluate the models on paragraph, line, character, and interleaved handwritten text generation, with results presented in Fig. 10. At the paragraph level, most models struggle with Chinese prompts. Only GPT-4o successfully generates coherent, clear paragraphs with accurately rendered text. Model performance improves as tasks simplify to line and character levels, particularly in the English scenario. For line generation, GPT-4o and Qwen-Image produce correct and clear sentences, followed by OmniGen2 with nearly perfect output. Other models only capture some keywords, while the rest of the text is poorly formatted. At the character level, most models can successfully generate the English letter “P”. For Chinese characters, however, only GPT-4o, Qwen-VLo-Preview, and Qwen-Image generate correct contents. For interleaved image-text generation, which requires generating a diagram with explanatory text, GPT-4o demonstrates superior capabilities. It accurately generates a hand-drawn physics diagram of the law of reflection, complete with clear and correctly placed labels. Other models either produce inaccurate diagrams or render the textual labels illegible.
4.2. Text Editing
We perform text editing on handwritten text images, with results shown in Fig. 11. At the page level, although GPT-4o achieved relatively optimal results, the output image is inadvertently cropped to a square shape and some content is lost. Other models perform even less satisfactorily. For example, Qwen-VLo-Preview misinterprets the prompt and inserts a stamp with incomplete text. Qwen-Image replaces the title with the target text but does not produce the embossed effect. Flux.1-Kontext-dev and OmniGen2 simply output the original image without any alteration. At paragraph and line levels, GPT-4o and Qwen-VLo-Preview successfully replace text but fail to maintain the original writing style and content, with issues such as content loss or errors, while other models fail entirely. The results reveal that most current models lack the capability for fine-grained editing. They often rely on a complete redrawing approach, which cannot retain the necessary style and background attributes of the original image. This highlights significant room for improvement in this task.
4.3. Handwritten Text Removal
In addition, we evaluate the handwritten text removal task, with results presented in Fig. 12. Two evaluation settings are adopted: erasing all text or erasing partial handwriting. Under the all-text removal scenario, given an image with simple structures, (i.e., comprising a single graphic, text, and background as shown in the first row), most models successfully erase all handwritten text within, except Qwen-Image, which produces two puzzling words. Conversely, when presented images with complex structures where handwritten text is embedded inside and interleaved with printed text, like the third row, most of them fail to precisely remove the handwriting. For partial removal, as shown in the second and fourth rows, models generally perform poorly. They often fail to precisely locate the target text, mistakenly remove other text that should be preserved, and alter the structure or appearance of the background. These results highlight the shortage of current generative models in removing handwritten text, underscoring the necessity of dedicated optimization.
5. Result of Scene Text Image
Figure 13. T2I generation results for scene text images. We test three language cases: English, Chinese, and mixed-language. Most models can handle English scene text generation. Yet, only GPT-4o and Qwen-VLo-preview handle Chinese text generation; Flux.1-Kontext-dev even delivers a totally unrelated result. For mixed-language, only Qwen-Image delivers a nearly perfect result, followed by GPT-4o and Qwen-VLo-Preview. Other models fail to handle this case, demonstrating chaotic, blurred, and nonsensical generated text. For complete prompts, please refer to our GitHub repository.
Figure 14. Text editing and text removal results for scene text images. In scene text editing, while GPT-4o successfully modifies the target text, it adds unexpected text (the first row) and alters the original aspect ratio. Other models exhibit inadequate text modification precision and substantial loss of textual and background texture fidelity. In scene text removal, models only fulfill the requirements in the first row case, while accidentally eliminating other details like the animal and auxiliary text, or failing to erase smaller texts.
Scene text, or text in the wild, refers to textual information that appears naturally in real-world environments, including vehicle license plates, street signs, and numerous other objects in daily surroundings. Unlike documents or handwriting, scene text features unconstrained font appearances and diverse backgrounds, such as curved text (Liu et al., 2020a), multiple text orientations (Zhou et al., 2017), perspective distortion (Shi et al., 2019), and low contrast and cluttered background (Lee and Osindero, 2016). Recent research has expanded beyond detection and recognition (Luo et al., 2019; Huang et al., 2022; Liu et al., 2023; Huang et al., 2024) to encompass generative tasks involving scene text (Liu et al., 2020b; Tuo et al., 2024; Chen et al., 2023; Peng et al., 2024), including scene text synthesis (T2I generation), scene text editing, and scene text removal. Therefore, we evaluate existing models across these three tasks to benchmark their capabilities in handling the complexity and diversity of text in real-world scenarios.
Figure 15. T2I generation results for images with naturally embedded text. In the first and second rows, the first three models perfectly follow the instructions and output applicable results, while other models either fail to generate specified text or synthesize blurred, unclear text. However, in the third and fourth rows, all models can not accurately generate ordered text like numbers on the ruler or letters on the keyboard. For complete prompts, please refer to our GitHub repository.
5.1. T2I Generation
Fig. 13 presents T2I generation result for scene text images. We evaluate three cases that feature English, Chinese, and mixed-language text to comprehensively demonstrate the models’ scene text generation capabilities across different languages. Most models can accurately generate English scene text, with only BAGEL and Janus-4o failing in this case. However, only GPT-4o, Qwen-VLo-Preview, and Qwen-Image can handle the Chinese scene text generation, while Flux.1-Kontext-dev produces a completely irrelevant image without any text, suggesting it may lack Chinese prompt understanding capability. For mixed languages, no model can accurately generate the text, though Qwen-Image produces a nearly perfect outcome (only “Bazaar” is wrongly rendered as “Baaar”). Other models mainly produce chaotic, blurred, or non-semantic content. In addition, we observe that Qwen-VLo-Preview struggles with other languages in the multilingual case, while GPT-4o and Qwen-Image deliver more balanced performance across diverse languages.
5.2. Scene Text Editing and Scene Text Removal
We also evaluate these models’ text editing and text removal capabilities on scene text images, as presented in Fig. 14. For text editing, only GPT-4o accurately modifies text according to instructions, but it still adds unexpected text (the first row) and does not preserve the original aspect ratio of the input image. The other models fail to modify the text and lose other textual or background texture details. For text removal, although some models can erase target text, many details remain problematic, such as changing background textures, failing to erase smaller text, or removing text and background that should be preserved. On both instruction-based editing tasks, Qwen-Image and Janus-4o demonstrate the worst performance: the former misunderstands instructions and edits the wrong text and the latter produces ruined images. These results indicate that scene text editing and removal remain unaddressed challenges for current generative models.
5.3. Naturally Embedded Text
Naturally embedded text refers to text seamlessly incorporated into physical elements (e.g., smartwatch, keyboard) rather than artificially overlaid. Due to its natural integrity, we define them as a type of scene text with different materials and integration methods. Hence, we specifically evaluate existing models’ T2I generation and text editing capabilities on naturally embedded text.
5.3.1. T2I Generation
We demonstrate the models’ T2I generation capabilities for naturally embedded text in Fig. 15. As observed in the first and second rows, the first three models perform excellently when generating electronic display fonts, and other models either fail to generate specified text or synthesize blurred, unclear text. However, when asked to generate items with ordered text, such as rulers or keyboards, all models struggle significantly. For instance, the rulers often show inconsistent or illogical measurement intervals and values, while the keyboards frequently display scrambled or repetitive characters instead of the standard QWERTY layout. This may be attributed to their lack of world knowledge or the insufficiency in synergizing their innate knowledge and generation performance, particularly in cases requiring precise spatial arrangement and sequential ordering of text elements.
Figure 16. Text editing results for images with naturally embedded text. GPT-4o successfully edits most texts but occasionally fails to preserve some original text or pixel texture details (the second and third rows). Other models primarily struggle to modify small text and directly output vanilla images or irrelevant results.
5.3.2. Text Editing
Furthermore, we evaluate text editing on naturally embedded text, with results shown in Fig. 16. We mainly instruct models to modify small text in images. Among all tested models, GPT-4o performs best, though it fails when modifying Chinese text (third row) and falls short in maintaining pixel texture details. Other models primarily struggle to modify much smaller text (like changing “5 km” to “63 km”) and sometimes directly output the vanilla input unchanged (OmniGen2 and BAGEL). Qwen-Image even generates a human picture given a fridge panel input. The results indicate that small text editing remains a significant challenge for existing generative models, particularly in cross-language applications and visual consistency maintenance. This provides clear directions for future optimization of multimodal models.
6. Result of Artistic Text Image
Figure 17. T2I generation results for artistic images. Only Qwen-Image generates complete content, while other models fail to completely and accurately render text lines with artistic fonts. Also, all models fail to generate images of rare Chinese characters. For complete prompts, please refer to our GitHub repository.
Artistic text is stylized typography that incorporates creative graphical components or decorative fonts to achieve aesthetic appeal beyond standard readable formatting. It has found widespread applications from creating memorable brand logos and advertisements to designing eye-catching movie titles, game interfaces, and personalized merchandise (Bai et al., 2024). In the recent decade, numerous synthesis methods (Gao et al., 2019; Wang et al., 2023; Li et al., 2023b; Tanveer et al., 2023) have been proposed to assist automatic glyph design, demonstrating remarkable font fidelity and diversity. Therefore, we incorporate this task into our evaluation, testing whether current generative models can fulfill this interesting, creative, and practical task.
Figure 18. Text editing and style transfer results for artistic images. Most models perform well on English script editing and style transfer. However, for Chinese text, only GPT-4o can handle editing and style transfer, though it still inadvertently crops the image. This indicates that current models’ multilingual text modification abilities require significant improvement beyond English.
6.1. T2I Generation
We first assess the T2I generation task, with results presented in Fig. 17. For line-level text generation, Qwen-Image is the only model that generates correct content in simplified Chinese and English scenarios. While other models produce visually pleasing artistic fonts, they suffer from content missing. In addition, Additionally, when tasked with generating complex, rare Chinese characters, they consistently fail to meet requirements. Flux.1-Kontext-dev performs worst, which generates human and photo frame images instead of text images. For character-level generation, all models can generate English single-character. However, they again fail to generate rare Chinese characters.
6.2. Text Editing and Style Transfer
Subsequently, we evaluate the text editing and style transfer abilities of current generative models. Results are shown in Fig. 18. Here, style transfer looks similar to text editing due to their shared objectives of modifying the source text to the target. Hence, it is necessary to clarify the definitions of these two tasks: text editing requires the models to modify the text and keep other elements of the image unchanged, while style transfer aims at generating a new image with specific text content according to the source image’s text style, regardless of other elements in the vanilla images.
We evaluate text editing under two scenarios: editing the text without and with style modifications. In the first and second rows, we instruct models to modify text while preserving the original style. Most models demonstrate strong performance on English text. For Chinese text, only GPT-4o and Qwen-VLo-Preview can perform the task, although both introduce artifacts: GPT-4o alters the background to black, while Qwen-VLo-Preview introduces unnecessary decorative elements. In the third row, most models successfully modify both text content and writing style, while preserving background elements like fireworks and cityscapes. Notably, Qwen-Image adds the new text without replacing the original, and models like GPT-4o and Janus-4o can turn the original rectangular images into squares, which is an unexpected minor issue.
For style transfer, the source style is well transferred to the target content in English cases. However, similar to text editing, most models can not handle style transfer of Chinese text. Collectively, these findings indicate that current models can handle English artistic text editing and style transfer effectively but require further improvements for Chinese script processing.
7. Result of Complex & Layout-Rich Text Image
Complex and layout-rich text images, such as slides and posters, blend text, graphics, and intricate layouts for both visual aesthetics and effective communication, which are widely used in marketing, education, and entertainment. While automated generation of these images grows increasingly popular, it requires understanding both textual semantics and design principles. For instance, slides require text seamlessly aligned with graphics to emphasize key points without clutter (Fu et al., 2022; Bandyopadhyay et al., 2024; Zheng et al., 2025), while posters demand bold typography and striking visuals to capture attention (Pang et al., 2025; Chen et al., 2025c; Peng et al., 2025; Gao et al., 2025). This drives us to include this type of text image in the evaluation.
Figure 19. T2I generation and text editing results for slide images. GPT-4o and Qwen-Image outperform other models in instruction-following and generation quality, meeting most requirements. However, other models generate unknown languages when English/Chinese is requested, also yielding square images despite instructions for rectangular ones. Regarding text editing, all models struggle with dense slide text. This proves that locating and modifying a small portion of text remains a great challenge for current generative models.
Figure 20. T2I generation and text editing results for poster images. Most models can handle the generation of English posters but fall short in generating Chinese posters, primarily attributed to their lack of Chinese training data. Qwen-Image delivers the only successful Chinese poster generation case. Similar to Fig. 19, models largely fail at accurate text editing, and even when edits succeed, original details are often altered.
Figure 21. Layout-aware text generation results for CLR text. For English cases, only OmniGen2 and GPT-4o add correct text in the proper position without occluding the main component, also preserving other details. For Chinese cases, only GPT-4o succeeds, while other models struggle to comprehend or generate Chinese text.
7.1. Slide Image
7.1.1. T2I Generation
We first investigate the slide generation abilities of existing models, with results presented in the top frame of Fig. 19. As observed, GPT-4o and Qwen-Image are the top two performers, which demonstrate stronger instruction-following abilities and text rendering abilities than other models, despite some content errors. In contrast, other models generate illegible text or even totally unrelated images. For example, Qwen-VLo-Preview produces an image of handwritten text but a slide. Similarly, Flux.1-Kontext-dev fails to interpret Chinese instructions, instead generating an unrelated image of a young girl. These results highlight the challenges generative models face in producing complex slide content.
7.1.2. Text Editing
We also evaluate their text editing capabilities of slide images, with results illustrated in the bottom part of Fig. 19. The overall performance is unsatisfactory. GPT-4o partially applies edits but disrupts the original layout and modifies accompanying images. Flux.1-Kontext-dev similarly fails to locate the source text and preserve the layout. Qwen-Image modifies the wrong text. Qwen-VLo-Preview and OmniGen2 alter backgrounds and text extensively, yielding unreadable results. BAGEL fails to respond to the instructions and leaves the image unchanged, while Janus-4o even produces chaotic and nonsensical content. Given rich and dense text in the slide, models struggle to locate a small portion of text while also preserving other elements.
7.2. Poster Image
7.2.1. T2I Generation
Similar to slide images, poster stands as an important type of visually rich image worth studying. Results of poster generation are shown in the top frame of Fig. 20. GPT-4o delivers the most consistent results, correctly and clearly rendering all English text, despite failing to render some small Chinese text (the third row). Qwen-Image is the only one that correctly generates small Chinese text, while the second generated poster somewhat lacks visual appeal. Other models demonstrate certain abilities in text generation, with frequent errors or omissions, especially in Chinese. Flux.1-Kontext-dev even produces a gate image instead of a poster. Models like OmniGen2, BAGEL, and Janus-4o frequently produce unreadable or nonsensical text, rendering their outputs unsuitable for practical use.
7.2.2. Text Editing
Results of text editing illustrated at the bottom of Fig. 20, GPT-4o correctly modifies partial requested content but also unexpectedly alters unrelated design elements. Qwen-VLo-Preview partially handles Chinese text editing but generates incomplete or incorrect modifications, while Qwen-Image misplaces text and produces unreadable content. Remaining models either fail to edit Chinese text or produce outputs with unreadable results. Except for GPT-4o, none successfully locate and edit the correct English text, and all introduce unintended changes to other elements. This indicates significant room for improvement on this task, particularly for open-source models.
7.3. Layout-aware Text Generation
Layout-aware text generation, also known as content-aware layout generation (Hsu et al., 2023; Horita et al., 2024; Seol et al., 2025; Guo et al., 2025), requires placing text naturally within images without obliging the original graphical component. Results are shown in Fig. 21. GPT-4o performs best, which correctly adds text without compromising the original layouts in most cases, although it may inadvertently alter the image (the first row). Qwen-VLo-Preview preserves layout but often renders text incompletely. Other models primarily produce erroneous or unreadable text, or sometimes leave the images unchanged. Specifically, Flux.1-Kontext-dev redundantly generates two lines of “Camera is good”, Qwen-Image hallucinates human hands (the last row), and OmniGen2 generates English text given a request for Chinese (the last row). These results highlight the challenge of this task for these models.
8. Conclusion
In this paper, we comprehensively evaluate the image generation/editing capabilities of several SOTA generative models across a spectrum of OCR generative tasks. We select 33 common OCR tasks and categorize them into five main groups based on text characteristics. Seven cutting-edge generative models are then evaluated, including two closed-source models, i.e., GPT-4o (OpenAI, 2025) and Qwen-VLo-Preview (Team, 2025) (unified generative models), and five open-source models, i.e., Qwen-Image (Wu et al., 2025b), Flux.1-Kontext-dev (Labs et al., 2025), and OmniGen2 (Wu et al., 2025d) (specialized generation models), as well as BAGEL (Deng et al., 2025) and Janus-4o (Chen et al., 2025b) (unified generative models). Our evaluation reveals that text image generation and editing remain significant challenges for current models, in which they suffer from inaccurate text locating, poor structural preservation, blurred or illegible character generation, etc. By providing the first dedicated evaluation of OCR-focused image generation, our work not only exposes the limitations of current generative models but also points out potential directions for future development. We envision that high-quality text image generation and editing could be internalized into general-domain generative models as foundational skills, and we hope our analysis inspires the community to advance toward this goal.